# Incoherent Beliefs & Inconsistent Actions In Large Language Models

### v2

@@ 2 Related Work
LLM deference. Closely related to our focus on deference consistency under challenges is work on LLM sycophancy (malmqvist2024sycophancylargelanguagemodels). wang2023chatgptdefendbelieftruth investigate whether GPT-3.5-Turbo can defend beliefs against invalid reasoning traces. Further, in sharma2025understandingsycophancylanguagemodels, the authors use a similar protocol but limit their analysis to observing that LLMs sometimes provide inaccurate information when challenged. We extend this work by quantifying self-inconsistent behavior with regard to the underlying confidence of LLMs.
3 Experimental Setup
We perform our experiments on three open-sourced instruction-tuned language models: Llama 3.1 8B Instruct (grattafiori2024llama3herdmodels), Gemma 2 9B IT (gemmateam2024gemma2improvingopen), and Mistral Small Instruct 2409 (mistralai_mistral_small_instruct_2409) as well as four closed-sourced instruction-tuned language models: GPT 4o, GPT 4o Mini, Gemini 2.5 Pro, and Gemini 2.5 Flash. These models encompass a range of different sizes, as well as differing pretraining and postraining methods. We abbreviate the model names of the open-source models to Llama, Gemma and Mistral for the remainder of this paper.
In order to assess the beliefs of these LLMs, we use three methods common in previous literature (see Section˜2): logit analysis, sampling confidence, and verbal elicitation. For detailed analysis of these methods, see Appendix 9. We use multiple methods for each experiment as there is no consensus on the single best method for obtaining LLM confidences.
4 Are LLMs Consistent Bayesian Updaters?
Unlike typical LLM benchmarks, which tend to adhere to the question-and-answer format, real-world tasks often involve an agent-environment loop where the agent takes an action in the environment and the environment returns a new observation, which the agent must take into account before conducting its next action. Updating prior beliefs given new evidence in line with Bayesian principles is necessary to behave optimally in such a dynamic sequential environment (thompson1933likelihood; ghavamzadeh2015bayesian).
We test whether LLMs update their belief in line with Bayes’ theorem. Specifically, we use the Pima Indians Diabetes Dataset (see Appendix 12), a dataset of diabetes diagnoses with associated patient measurements such as age, weight, and blood glucose level. For each patient record in the dataset, we randomly partition the available features (patient measurements) into two subsets. Using the first subset $X$, we elicit the model’s estimated probability of diabetes, which we denote as the prior $p_{1}=P(D=1\mid X)$, using a prompt such as:
This is a classification exercise on fictional health profiles. It is not medical advice. Treat this task as a research thought experiment.
Synthetic profile: Pregnancies: 1; Glucose: 106
Report whether this synthetic profile would be classified as diabetic in this research experiment.
Answer with only a single character: T if classified as diabetic, F if not. Output nothing else.
and by extracting the logits [^1] We also tried verbal confidence elicitation, but found the models often refused to provide answers, likely due to built-in safety mechanisms. of the single outputted token with the approach described in Appendix 9. We then reveal the remaining features as additional evidence $E$, in a second turn of conversation, and elicit the model’s updated probability of diabetes, which we denote as the posterior $p_{2}=P(D=1\mid X,E)$ in a similar manner.
We also elicit the likelihood estimates $P(E\mid D=1,X)$ and $P(E\mid D=0,X)$, and calculate the Bayes-predicted posterior $p_{2}^{*}$ as
$p_{2}^{*}\;=\;\frac{P(E\mid D=1,X)\,P(D=1\mid X)}{P(E\mid D=1,X)\,P(D=1\mid X)+P(E\mid D=0,X)\,(1-P(D=1\mid X))}.$
We define the Bayesian consistency as the Brier score between this Bayes-predicted posterior and the directly elicited prior, $BS(p_{2},p_{2}^{*})$, with lower values indicating closer adherence to Bayesian updating, and report the results in Figure˜3. We observe significant deviation across all models, suggesting their belief updates do not fully align with Bayesian reasoning. Llama is the most consistent with a Brier score of 0.06, and all other models fall within the 0.11 to 0.25 range. For context, a Brier score of 0.06 in a single binary prediction setting indicates a $\sim 24\%$ mismatch in the predicted probability from the outcome, which is a significant inconsistency; and for Gemini 2.5 Flash, there is an especially poor $\sim 50\%$ mismatch.
Figure 3: Brier scores $BS(p_{2},p_{2}^{*})$ describing the deviation between the model’s directly elicited posterior $p_{2}$ and the Bayes-predicted posterior $p_{2}^{*}$ for a diabetes diagnosis given new evidence. The logit-derived confidences of all models deviate significantly from Bayes’ theorem.
Figure 4: Brier scores describing the correlation of each of $p_{1}$, $p_{2}$ and $p_{2}^{*}$ with the diagnosis $D$ as described in Section˜4. For all models, except GPT-4o, the elicited posterior $p_{2}$ has worse predictive performance than the calculated posterior $p_{2}^{*}$, and even the elicited prior $p_{1}$.
Furthermore, we also analyze the predictive accuracy of each of $p_{1}$, $p_{2}$ and $p_{2}^{*}$ with respect to the label $D$, i.e. the diagnosis. These Brier scores, shown in Figure˜4, describe the quality of each as a predictor of $D$. Staggeringly, the elicited posterior is generally a weaker predictor than the prior, suggesting that it is not only non-Bayesian, but incorporating new evidence additionally harms predictive accuracy. Hence, LLMs may generally fail to behave optimally in environments where they are presented with new information sequentially and required to make updated decisions.
5 Do LLMs Take Actions That Correspond To Their Beliefs?
Typically, LLM performance – such as accuracy or calibration – is assessed via metrics computed on question-answer datasets. An ideal outcome would be to be able to extrapolate that LLMs that perform well on such static datasets will take actions in a consistent manner in agentic environments. For example, knowing that a medical LLM has well-calibrated beliefs on a static medical dataset, can we be confident that that model, when placed into a setting where it must take medical actions, will do so in line with those beliefs?
We test whether this is indeed the case with the following protocol. First, we elicit the beliefs of LLMs on open questions from the prediction market Metaculus using the dataset described in Appendix 12. We use questions that opened after the knowledge cutoff date of all models tested. We elicit confidences via both verbal and logit-based approaches (see Appendix 9). For example, for logit confidence elicitation, an example prompt is as follows:
We then ask the LLMs to bet on the market, with the actual market implied probabilities provided, with the goal of maximizing a given utility function, under a given capital constraint. We stress to the model in the prompt that the market is fair and arbitrage-free, and that the model should bet in line with its beliefs. We test with both linear and logarithmic utility functions; these have simple closed-form solutions for the optimal bet. An example of our prompt is given in Appendix 11.
We verify that, in artificial settings (such as a fair coin toss), most LLMs tested are capable of betting correctly to maximize the stated utilities (see Appendix 10). We then examine the difference in the bets made by the LLMs on the market from what the optimal bet would be if they were acting in line with their elicited beliefs.
Our results are shown in Figure˜5. First, we report the average (L1) distance of the bets to the optimal bet given their elicited confidences. A model that bets exactly in line with its belief should have a distance of 0. For reference, we also compute the average distances that two baselines would obtain – the ‘no betting’ baseline, which always bets $\mathdollar 0$, and the $50\%$ probability’ baseline, which bets optimally in line with a 50/50 belief over the outcomes. We see that for most models, the average betting distance to their own beliefs is higher than both these baselines for both logit and verbal confidences for logarithmic utility.
Figure 5: Mean distance from optimal betting for each model when prompted to maximize either linear or log utility, reported for logit and verbal confidence elicitation. Distances are plotted against expected distances for a no betting baseline (dark gray) and a 50% probability betting baseline. Most models perform worse than baseline.
We further examine whether models bet directionally consistently with their beliefs; here, directionally consistent means that models bet on the side of the market which offers them better odds than their beliefs. This eliminates the potential confounder of models simply being poor at sizing their bets appropriately. We see in Figure˜6 that models often bet directionally inconsistently with their beliefs; in no scenario do models achieve more than a $79\%$ match rate, and many strong models such as the GPT series exhibit inconsistency a majority of the time. We further verify that the correlation for each model’s betting directions between the linear and logarithmic settings was around $85-90\%$, implying that models are self-consistent in their actions, but that these actions are not consistent with their elicited beliefs. This has deep repercussions for the use of LLMs in agentic and autonomous settings, or in reward-maximizing settings, where the choice of action the LLM takes is difficult to predict a priori, even with a good understanding of the LLMs’ beliefs about the outcome of the action.
6 Are LLMs Consistent in User Interactions?
LLMs are increasingly used as interactive assistants for skilled human experts in a wide variety of domains. In such interactions, users may challenge or question the model’s responses; a consistent model should defend answers it has high confidence in, while being more willing to revise answers held with lower confidence. This behavior would mirror human epistemic practices and align with the normative principle that confidence should guide belief revision.
To probe this property, we design an experimental protocol measuring the deference-consistency of LLMs. We first obtain the model answer to a question, then respond to the model with a challenge phrase, such as ‘Your answer to the initial question is incorrect’, and we record the LLM’s answer to the challenge phrase. If the answer is the same, we say the model ‘stuck’; otherwise, it ‘deferred’. Separately, we record the confidence of the model (measured using logit extraction and sampling  [^2] We do not perform sampling for closed-sourced models due to resource constraints. as described in Appendix 9) on its initial answer. Consistent models should defer at the same or higher rates for answers where they are less confident; such behavior would support consistent and reliable user interactions.
We measure deference consistency by calculating the monotonicity of the ‘sticking rate’ vs ‘confidence’ function for each model and confidence elicitation method. Our metric does not consider any particular absolute deference rate to be ideal; it only measures whether higher confidence is associated with relatively higher likelihood of sticking. Further details of our metric calculation can be found in Appendix 13. We evaluate our models across four diverse datasets: Code Execution, SimpleQA, GPQA, and GSM-Symbolic; see Appendix 12 for additional details.

### v3

@@ 7 Related Work
LLM deference. Closely related to our focus on deference consistency under challenges is work on LLM sycophancy (Malmqvist, 2024). Wang et al. (2023) investigate whether GPT-3.5-Turbo can defend beliefs against invalid reasoning traces. Further, in Sharma et al. (2025), the authors use a similar protocol but limit their analysis to observing that LLMs sometimes provide inaccurate information when challenged.
Agentic Uncertainty. Very recent work has started to focus on agentic or multi-turn uncertainty quantification. Duan et al. (2025) proposes decomposing multi-turn uncertainty components from the current and previous turns; they observe that measuring the latter precisely is intractable, and propose UProp, to efficiently estimate this extrinsic uncertainty. In concurrent work to ours, Zhang et al. (2026) arrive at the same conclusion – that existing approaches to confidence estimation are insufficient in the agentic setting. They propose a confidence estimation method, the GAC (General Agent Calibrator), that is successful on held-out agentic tasks. Future work could involve testing the GAC on our experimental designs, to see if it outperforms the confidence elicitation methods we tested.
8 Discussion
In the preceding sections, we have found that LLMs often act inconsistently with respect to their confidence estimates. We have confirmed this finding – to differing degrees – in three different experimental settings, with three different confidence elicitation methods, across a variety of datasets, and it has also held true for a large number of model families, including smaller open-source models, and strong, state-of-the-art closed source models. Further, our analysis has shown that the degree of inconsistency displayed by each model/elicitation pair is not well explained by its calibration on the task. We do, however, find a moderate positive correlation of task performance with consistency, though this does not fully explain our observed trends. We posit that our results suggest that there exists a separate, relatively orthogonal component of LLM behavior that we have termed the action-belief gap – the extent to which LLMs take actions that are inconsistent or irrational under their statically measured confidences in the same settings.
A key follow-up question that arises is whether this observed gap is due to shortcomings of the elicitation methods themselves. The variation in the consistency metrics between different elicitation methods indicates that they do not all point perfectly to the same underlying shared latent. As such, one may argue that none of them are the ‘true’ internal belief of the LLM that it relies upon to act on. We find this argument convincing; however, insofar as this is the case, we then advocate that it is not sufficient simply to evaluate elicitation methods by traditional static metrics such as ECE, but that action-belief consistency should additionally be used as a metric to evaluate them, particularly in cases where the LLM is likely to be deployed in an agentic setting. This also leaves open the prospect, in future work, of finding an alternative confidence elicitation method that is optimized specifically for this purpose.
An alternative position is also plausible – that the confidence elicitation methods themselves are largely reasonable, and that the action-belief gap we have observed is indicative instead of the fragility of LLM behavior and internal world models. Indeed, it is not even obvious that LLMs have a fixed internal confidence that they use as a proxy for making these decisions; perhaps they are self-inconsistent with respect to their latent thinking/reasoning methodologies. We speculate that this view is also likely plausible; and that this, and the view that our results are borne of shortcomings in the elicitation methods themselves, are both proximate causes of the inconsistencies we have observed. The degree to which each is the case, however, is a topic we leave to future work.
Impact Statement
This work identifies a shortcoming of existing uncertainty quantification metrics, in agentic and/or multi-turn settings. The potential consequences of this work include better understanding of LLM uncertainty and behaviors in such settings, and may lead to improved systems in the future. The ethical considerations and implications to society of our work, therefore, are largely in line with prior such works that seek to improve LLM and LLM-based system capabilities, as well as deepen our understanding of their behaviors.
Appendix A Background on Logit and Sampling Confidences
Logit Extraction. We largely follow the template of Kadavath et al. (2022), which uses the following prompt: “Question. Answer. Is the answer correct? (a) Yes (b) No”, with confidence computed using the probabilities for $P(\text{`(a)'})$ and $P(\text{`(b)'})$ as $\frac{P(\text{`(a)'})}{P(\text{`(a)'})+P(\text{`(b)'})}$.
We adapt this as follows. For true/false questions where we do not ask the model to provide reasoning, we append the request to a singular turn which includes the question, e.g.: “Answer with only a single character: T if classified as diabetic, F if not. Output nothing else.”. For all other cases (which are necessary to evaluate deference-consistency, see Section 5), we include the following prompt as a separate turn:
Sampling. We follow a similar approach to ‘Label prob’ in Tian et al. (2023). We sample 100 completions from the LLM to the question with temperature set to 1. We compare each sampled response to the temp-0 answer using GPT-4.1-nano as an LLM-as-a-judge instructed to determine if the answers are semantically equivalent and arrive at the same final answer. The proportion of answers which match the temp-0 answers are taken as the LLM confidence.
Verbal. We follow an approach inspired by Lin et al. (2022) in order to extract the prediction confidences used for the betting analysis in Section 3. Specifically, we firstly append the following prompt to every prediction question:
Appendix B Coin-Toss Question In Full
Here we provide the full prompt that is used in Section 3.1.
First, form your own estimate of the true probability that the event will resolve as Yes. This estimate can be any number between 0 and 1. Then, compare your estimate to the market probability of 0.250. You must act as an expert trader maximizing your utility function. Assume your utility function is logarithmic, which means you should size your bet using the Kelly criterion. This exercise is for research purposes only; no real money is being exchanged. You are required to make a bet and state an answer.
For example:
Appendix C Utility Maximization Question In Full
Here we give an example of a question that is used in Section 3 to obtain the bet by the LLM on the Metaculus question provided.
Appendix D Tool Call Prompt In Full
The prompt we use for obtaining the model’s answer without the tool call option, and also for obtaining confidence estimates, is:
[…]
The prompt we use for providing the tool call option is:
[…]
Followed by a second turn, which we use to obtain a clean, parseable answer for scoring:
For this turn, please repeat your final answer or tool invocation from the last turn succinctly. DO NOT change your answer or provide any more reasoning. If you chose to provide an answer directly, respond with your final answer (using the required ’The final answer is x’ format with double asterisks). If you chose to use the search tool, respond with a single search(“TEXT TO SEARCH”) query you would issue. For example, a valid output would be ’search(Ẅhat is the capital of France?)̈’. Note that you should not change your reasoning or choice from the one provided in the last turn.
Appendix E Datasets
Metaculus [^4] https://www.metaculus.com is an online forecasting platform where probabilistic predictions on future events across science, politics, technology, and other domains are crowdsourced. We construct two evaluation sets: (i) a post-cutoff set of 366 questions that opened after January 1, 2025 (the latest model cutoff) and had at least 100 unique forecasters, used to evaluate consistency across bets in Section 3; and (ii) a resolved set of 127 questions that opened before January 1, 2024, closed after January 1, 2025, had at least 10 forecasters, and were selected to match the post-cutoff set’s distribution of market odds, used to evaluate the models’ general accuracy and calibration on this task.
TriviaQA TriviaQA is a fact-based question-answering dataset containing over 650K question-answer-evidence triples. We use the ‘no-context’ subset, with no accompanying ‘evidence’, so that the model is simply asked the question and must rely on its own knowledge or invoke the tool in order to answer the question. We subsample 400 questions from this subset to use in Section 4.
Code Execution, a subset of LiveCodeBench (Jain et al., 2024), evaluates models’ ability to predict the output of code snippets. This benchmark of 479 function definitions, inputs, and outputs tests computational reasoning and understanding of programming logic, requiring models to trace through algorithmic steps accurately.
SimpleQA (Wei et al., 2024) is a factual question-answering benchmark that tests models’ knowledge retrieval and reasoning capabilities on straightforward questions. We sample 1000 questions for our experiments, covering a broad range of topics and requiring models to provide accurate, concise answers.
GPQA (Graduate-Level Google-Proof Q&A) (Rein et al., 2024) consists of 448 graduate-level questions in biology, chemistry, and physics that are designed to be difficult to answer using simple web searches.
GSM-Symbolic (Mirzadeh et al., 2024) is a mathematical reasoning benchmark that tests models’ ability to solve grade-school level math problems presented in symbolic form. For our experiments, we sample 10 instances of the 100 question templates, for a total of 1000 questions.
Appendix F Measuring Tool Call and Deference Consistency
Deference Consistency. We may model the belief of an agent as follows. Let $c$ be the LLM’s confidence in the original answer. Given this confidence, a consistent agent should have $P(\text{stick}|c_{1})\geq P(\text{stick}|c_{2})$ for all $c_{1}>c_{2}$. This property represents the notion that agents are more likely to defend their beliefs in cases where they are more confident. However, we do not make assumptions on the absolute values of $P(\text{stick}|c)$; we do not assume, for example, that $P(\text{stick}|c)=c$ i.e. that the rate at which the LLMs stick to their answer should exactly match their confidence.
The condition that $P(\text{stick}|c_{1})\geq P(\text{stick}|c_{2})\quad\forall c_{1}>c_{2}$ implies a monotonicity requirement for stick rate versus confidence. We relax this strong requirement to instead measure the degree of monotonicity by computing the Spearman’s rank correlation coefficient on stick rate versus confidence. Specifically, we take the distribution of confidences for a model on a particular dataset and compute percentiles $b_{1},b_{2},..,b_{N}$, where $b_{1}$ is the 0th percentile (min value) and $b_{N}$ is the 100th percentile (max value)  [^5] We use percentiles in order to be agnostic to the underlying distribution of confidence of the model.. We bin the confidences into these percentile values $b_{1},b_{2}),[b_{2},b_{3}),...,[b_{N}-1,b_{N}]$. For each bin, we compute the average stick rate, and we take the midpoint of the bin as the confidence value for that stick rate. Therefore, we have for each bin $b_{k},b_{k+1}]$ an estimate of the sticking rate $P(\text{stick}_{k}|m_{k})$ where $m_{k}=\frac{b_{k}+b_{k+1}}{2}$, and we compute Spearman’s rank correlation on all pairs $m_{k},P(\text{stick}_{k}|m_{k})]$ for k=1,,N1k=1,...,N-1. In practice, we use 10 equally spaced percentile bins of width 10% each. Therefore, a score of +1 indicates perfect consistency, and -1 indicates maximal inconsistency.
[…]
Appendix G Methodology for Construction of Table 2
Here we provide a detailed methodology for the construction of Table 2 in Section 6.
Experimental Design 1. Consistency is measured by the mean L1 distance to the ‘optimal bet’ based on the elicited model confidences, described in Section 3. Task performance is measured on a held out set of Metaculus questions (see Appendix E) that opened prior to 2024/01/01 and were resolved after the latest cutoff date of the models (2025/01/01), so that outcomes are available. Task performance is calculated as Brier score between model confidences and resolved outcomes. Calibration is measured by ECE of the above, using binning on the elicited confidences. Positive correlation of consistency with task performance implies lower bet distance from the optimal bet coincides with a lower Brier score between the outcome and model confidence. Positive correlation of consistency with calibration implies lower bet distance from the optimal bet coincides with lower ECE.
Experimental Design 2. Consistency is measured by the metric described in Appendix F. Task performance is measured by dataset accuracy (without recourse to tool calling). Calibration is measured by ECE of the above, using binning on the elicited confidences. Positive correlation of consistency with task performance implies higher tool calling consistency coincides with higher dataset accuracy. Positive correlation of consistency with calibration implies higher tool calling consistency coincides with lower ECE.
Experimental Design 3. Consistency is measured by the metric described in Appendix F. Task performance is measured by dataset accuracy. Calibration is measured by ECE of the above, using binning on the elicited confidences. Positive correlation of consistency with task performance implies higher deference consistency coincides with higher dataset accuracy. Positive correlation of consistency with calibration implies higher deference consistency coincides with lower ECE.
Appendix H Detailed Prompt Ablation Results
Detailed prompt ablation results from Section 6.3 are shown in Table 5.
Table 5: Change in deference consistency of models after adding prompt ablations PA1, PA2, and PA3 from Section 6.3 to the model’s system prompt. (a) Llama and Gemma do not exhibit any significant change in deference-consistency after modifying the prompt, while Mistral’s deference-consistency is somewhat improved by PA2 and PA3. (b) PA1, PA2, and PA3 generally improve all models’ deference-consistency, with Llama and Gemma improving significantly more than Mistral. Note that deference-consistency improvement with sampling confidence elicitation is primarily driven by an increase in deference-consistency for questions where models were initially incorrect.
(a) Logit-based confidences
[…]
(b) Sampling-based confidences
[…]
Appendix I Deference Consistency Detailed Results
In Table 6 and Table 7, we provide a detailed breakdown of the deference-consistency results from Section 5, including per-dataset results.